OpenStudy (anonymous):

How do you find the slope of(-1/2,-2) & (-3/2,1)?

OpenStudy (anonymous):

what you wanna do is use this slope formula: m=(y2-y1)/(x2-x1), where m is the slope

OpenStudy (anonymous):

I know how to do that but what do we do with the fraction to get it to be one whole number?

OpenStudy (anonymous):

@vanpai you just have to work with the fractions. no one says a slope is a whole number.

OpenStudy (anonymous):

if you use the formula you will get: m=((1)-(-2))/((-3/2)-(-1/2)) m=(3)/(-2/2)=(3)/(-1)=-3

OpenStudy (anonymous):

Thank you